OpenStudy (anonymous):

HELLPPPP!!!! PLLLEEZZ!! :)) Which of the following is the solution to the equation sqrt5^8n = 125^(n + 5) ? n=5,n=-5,n=15,n=-15?

OpenStudy (helder_edwin):

\[ \large \sqrt{5}^{8n}=125^{n+5} \] ??

OpenStudy (anonymous):

u can see here http://www.wolframalpha.com/input/?i=sqrt5%5E8n+%3D+125%5E%28n+%2B+5%29

OpenStudy (anonymous):

@helder_edwin yah thats right ;)

OpenStudy (helder_edwin):

Ok first \[ \large \sqrt{5}^{8n}=(5^{1/2})^{8n}=5^{1/2\cdot8n}=5^{4n} \] OK?

OpenStudy (anonymous):

yah!

OpenStudy (helder_edwin):

then \[ \large 125^{n+5}=(5^3)^{n+5}=5^{3(n+5)}=5^{3n+15} \]

OpenStudy (helder_edwin):

so \[ \large \sqrt{5}^{8n}=12^{n+5} \] becomes \[ \large 5^{4n}=5^{3n+15} \]

OpenStudy (helder_edwin):

agree?

OpenStudy (anonymous):

ok... but where did 12 come from?

OpenStudy (helder_edwin):

sorrry 125 not 12

OpenStudy (anonymous):

ok!! yah i understand so far! ;)

OpenStudy (helder_edwin):

now we have this \[ \large A^x=A^y\qquad\Rightarrow\qquad x=y \]

OpenStudy (helder_edwin):

this means that \[ \large 5^{4n}=5^{3n+15} \] becomes \[ \large 4n=3n+15 \]

OpenStudy (helder_edwin):

u can finish this right?

OpenStudy (anonymous):

so its right?